C=q^3+1.5q^2+9q......



Find AC .....



MC



AVC



AFC..?




1
Expert's answer
2022-03-29T12:10:21-0400

"AC = \frac{TC}{Q}"


"AC = \frac{{q^{3}+1.5q^{2}+9q}}{q}"


"AC = q^{2}+1.5q+9"


"MC = \intop" "\frac{dTC}{dQ}"


"MC = 3q^2 + 3q + 9"


"AVC = \frac{TVC}{Q}"


In this case; "AVC = AC" because there is no fixed cost

"AVC=q 2\n +1.5q+9"


"AFC = \frac{TFC}{Q}"

In this case; "TFC = 0; thus AFC = 0"